1. Using the definition C=Q/V, show that for a SERIES connection of capacitors: 1/C = 1/C1 + 1/C2. What does this equation mean? What happens to C as more and more capacitors are connected in series?
  2. Using the definition C=Q/V, show that for a PARALLEL connection of capacitors: C= C1+C2. What does this equation mean? What happens to C as more and more capacitors are connected in parallel?
1
Expert's answer
2021-05-14T10:19:12-0400

Gives

(a) series combination

First capacitor capacitance =c1

Second capacitor capacitance=c2


V=V1+V2

"\frac{Q}{C}=\frac{Q}{C_1}+\frac{Q}{C_2}"

"\frac{1}{C}=\frac{1}{C_1}+\frac{1}{C_2}"

"C<C_1,C<C_2"

(b)

Parrelled combination



Q=Q1+Q2

"CV=C_1V+C_2V"

"C=C_1+C_2"

"C>C_1\n,C>C_2"